OpenStudy (yoongilife):

I am willing to do the work. I want to understand! Topic: Chemical Thermodynamics I will post question once thread becomes open. Thanks~

OpenStudy (yoongilife):

You have the following equations: 2H2O (l) → 2H2 (g) + O2 (g), DELTAH = 572 kJ 2N2 (g) + 3H2 (g) → 2NH3 (g), DELTAH = -138 kJ 2NO2 (g) + 7H2 (g) → 2NH3 (g) + 4H2O (l), DELTAH = -1.18 X 10^3 kJ 2NO2 (g) → 2N2 (g) + 2O2 (g), DELTAH = ? What is the missing DELTAH? What I know: Honestly, chemical thermodynamics is not my strong point. Could we use Hess’s Law on this problem? I am willing to work, but I hope for a little bit of instructional guidance. Thank you!

OpenStudy (yoongilife):

@sweetburger I know I always call upon you when I am in distress over chemistry, but I believe you helped me with a chemical thermodynamic problem before. If you have time would you be interested in helping again? Thanks.

OpenStudy (sweetburger):

Honestly I am not quite sure how to go about this question but I have some ideas that may possibly correct. We want the Delta H of 2NO2 (g) → 2N2 (g) + 2O2 (g) I think we would assume that 2NO2 produces -1.18 X 10^3 kJ of energy, assume 2N2 produces -138kJ of energy, and assume that O2 produces -572kJ of energy. Then I would believe that you apply Hess's law. Dont forget to double -572kJ as the reaction in question is 2O2 not just O2.

OpenStudy (yoongilife):

True, that's how I kinda mapped it out as well. I think mostly I was looking for a second opinion.

OpenStudy (sweetburger):

Scratch the endothermic bit it actually might be exothermic.

OpenStudy (yoongilife):

As far as exothermic/endothermic goes I always want to verify, but I did the same thing. "Oh, it's endothermic!...Then again maybe not."

OpenStudy (yoongilife):

That being said, if you have time could I post one more question? I understand and respect if you are unavailable. Thank you for the help thus far!

OpenStudy (sweetburger):

I can still help out.

OpenStudy (yoongilife):

Thank you! Which of the following has a positive enthalpy without a temperature change? Cooling a solid. Condensing a gas. Evaporating a liquid. Evaporating a liquid is most definitely a temperature change. Cooling a liquid. Cooling a liquid is most definitely a temperature change. What I know: Enthaply includes the internal energy and the amount of energy required to make room for it by displacing its environment and establishing its volume and pressure. I'm not sure if I'm missing something, or if I'm looking at it in the wrong way.

OpenStudy (yoongilife):

Condensing a gas would increase collisions, right? As a result the internal energy would increase as well...

OpenStudy (sweetburger):

Cooling a solid has a temperature change and cooling a liquid has a temperature change. Actually Evaporating a liquid is a phase change therefore there is no temperature change as all the of energy that is inputted into this phase goes into breaking the bonds between the liquid molecules so they may become a gas.

OpenStudy (sweetburger):

It also has a positive enthalpy value as energy is taken in from the environment to go from liquid-->gas

OpenStudy (yoongilife):

Ah! Duh, how could I forget about phase changes! I feel slightly ashamed. Nevertheless thank you for bringing that to my attention. As always I appreciate your help and effort!

OpenStudy (sweetburger):

No problem always glad to help!